Use a suitable identity to get each of the following products.
 (2y + 5) (2y + 5)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(2y + 5) (2y + 5)=(2y + 5)^2
= 4y^2+ 20y + 25
Using (a+b)^2= a^2+ b^2 + 2ab

Thanks